# biburlnumpenalty

### v1

@@ Abstract
The localization game is a variant of the game of Cops and Robber in which the robber is invisible and moves between adjacent vertices, but the cops can probe any $k$ vertices of the graph to obtain the distance between probed vertices and the robber. The localization number of a graph is the minimum $k$ needed for cops to be able to locate the robber in finite time. The localization capture time is the number of rounds needed for cops to win.
@@ 1 Introduction
The Cops and Robber game on graphs has been widely studied; see for example [7]. The game is played on a graph by two players, one controlling the cops and the other the robber. Positions of cops and robber are known to both players at all times. The players are allowed to move to an adjacent vertex or stay in the same vertex, and players alternate taking moves. The goal of the cops is to capture the robber, i.e. at least one cop should occupy the same vertex as the robber. Robber’s goal is to evade capture forever. The minimum number of cops needed to ensure cops winning on a graph $G$ is the cop number $c(G)$ of $G$. In 2009, Bonato et al. [2] introduced the capture time of a graph as the minimum number of moves needed for $c(G)$ cops to win on $G$. The problem is well-understood for cop-win graphs, but only a handful of results are known for graphs with a larger cop number [15, 16, 12].
@@
The localization game is a variant of the Cops and Robber game introduced in 2012 by Seager [17]. The game has been further studied for example in [5, 4, 12, 9, 10, 6]. This game is also played on a graph by two players, one controlling the cops and the other the robber. However, the robber is invisible to the cops during the game, while the robber knows the positions of the cops. The players alternate taking moves, with the robber selecting the starting vertex first. The robber is allowed to stay on the same vertex or move to a neighbor. The cops are allowed to move to any set of vertices (not necessarily adjacent to their previous positions), and after they all move, each cop sends out a cop probe, which gives the distance between each cop to the robber. Thus after every move of the cops, they obtain a distance vector $(d(c_{1},r),\ldots,d(c_{k},r))$ where $c_{1},\ldots,c_{k}$ are positions of the cops and $r$ is the position of the robber. The goal of the cops is to determine the position of the robber after finitely many rounds. Robber’s goal is to evade capture forever. Additionally, we assume that the robber is omniscient, meaning that the robber is aware of all strategies of the cops. The localization number $\zeta(G)$ of $G$ is the minimum number of cops needed to ensure that the cops win the localization game on $G$. We only consider this game for connected graphs. The exact upper bound for the localization number is known for example for trees, outerplanar graphs [5], and some diameter 2 graphs [4], and is known to be unbounded for planar graphs [9].
@@
While specific variants of the problem have been studied before [17, 18, 11, 14, 13], in 2022, Behague et al. [1] introduced in full generality the analogous problem to the capture time for the Cops and Robber game for the localization game. The $k$-localization capture time $\operatorname{lcapt}_{k}(G)$ is the minimum number of rounds needed for $k\geq\zeta(G)$ cops to capture the robber on $G$ if the cops are minimizing the number of rounds and the robber is maximizing it. If $k=\zeta(G)$, we refer to the parameter as the localization capture time and denote it by $\operatorname{lcapt}(G)$. Recall that the game starts by the robber choosing an initial vertex. Then, round 1 begins. Each round consists of a move of the cops, after which they obtain a distance vector to the robber’s position (we also say that the cops probe the $k$ vertices they move to), which is followed by a move of the robber. The number of rounds needed for cops to determine the position of the robber is thus equal to the number of times the cops obtain a distance vector to the robber, i.e. the number of times the cops probe the vertices. Note that as $k$ increases, $\operatorname{lcapt}_{k}(G)$ decreases, and it equals 1 if $k$ is the metric dimension of $G$.
@@ 2 Trees
It is known that one cop wins the localization game on a tree $T$ if and only if $T$ does not contain a copy of $T_{3,3}$ ([18, Theorem 8]). As mentioned in [8] and explicitly proven in [3], it is known that if $T$ is a tree, then $\zeta(T)\leq 2$. It follows from [18] and [1] that the best known upper bound for the localization capture time for a tree $T$ on $n$ vertices is
@@
In this section we improve on these two upper bounds. As $\ell(T)\leq n-1$ for every tree on $n\geq 3$ vertices, and $\ell(T)\leq n-4$ for every tree on $n$ vertices that contains a copy of $T_{3,3}$, we get the following result from Theorems 2.2 and 2.3 which are proved below.
@@
If $T$ is a tree on $n$ vertices, then
$\operatorname{lcapt}(T)\leq\begin{cases}n-1;&\text{if }\zeta(T)=1,\\ \left\lfloor\frac{n}{2}\right\rfloor-3;&\text{if }\zeta(T)=2.\end{cases}$
We also mention below that the given upper bounds cannot be improved. In the proofs, the candidate set denotes the set of all possible locations of the robber on a graph after a move by the cops or the robber is completed.
@@
If $T$ is a tree that does not contain a copy of $T_{3,3}$, then $\operatorname{lcapt}(T)\leq\ell(T)$.
@@
Let $T$ be a tree which does not contain a copy of $T_{3,3}$, and root $T$ at any leaf $v_{0}$. We provide a strategy for the cop to construct a sequence of pairwise distinct vertices $v_{1},v_{2},\ldots,v_{a}$ of $T$, lying on levels $k_{1}<k_{2}<\cdots<k_{a}$ respectively, such that for each $1\leq i\leq a$, the candidate set will be contained in the set of descendants of $v_{i}$, immediately after a cop’s probe. In particular, after each application of one of the strategies described bellow, the cop will be able to either capture the robber, or identify the lowest positioned vertex $v_{i}$, on level $k_{i}>k_{i-1}$, such that the candidate set will be contained in the set of its descendants throughout the remainder of the game. Since the tree is finite, the cop will indeed capture the robber in a finite number of moves. Moreover, we prove that during the described strategy, the cop (in most cases) maintains the following property: the number of probes needed to eliminate a part of the tree from all the possible future candidate sets with respect to the outlined strategy is at most the number of leaves in that part of the tree. There are a few exceptions to this general rule which are dealt with on the spot.
Choose $v_{0}$ as the first probe. If the robber is not captured, then the candidate set is clearly contained in the set of descendants of $v_{0}$, so we may set $v_{1}=v_{0}$, or pick a lower positioned vertex if that is possible. Suppose that after some move, the cop has already constructed $v_{1},\ldots,v_{t}$. Since the robber must be located on a descendant of $v_{t}$, we may now consider the levels of these vertices with respect to $v_{t}$, that is, as if the tree was rooted at $v_{t}$, in order to simplify the notation. Recall that the vertex set consisting of a child of $v_{t}$ and all of its descendants is called a branch.
@@
Suppose now that there is at least one child of $v_{t}$ of degree 2. Again, if such children are not present, the cop simply omits this step. Otherwise, he consecutively probes the children of such vertices. Let $u$ be one such vertex that the cop probes, and let $d$ be the answer he gets after this probe. Immediately before the probe, the robber cannot be on any other level than k1,kk-1,k or $k+1$. We consider the following four cases. In all of them, if the cop ever realizes that the robber is on the same branch as him and on level 3 or lower, we may simply set $v_{t+1}$ to be $u$ or, if possible, some other vertex on the branch on a lower level. Where the number of probes is clearly at most the number of leaves in the eliminated part of the tree, we do not explicitly mention it.
@@
Notice that in each case, the cop has eliminated some branches as the robber’s possible whereabouts, using at most as many moves, as there are leaves on the discarded branches. Hence, using the described strategy, the cop extends his sequence of vertices, until the robber is captured, needing at most $\ell(T)$ moves to do so. ∎
Note that this upper bound cannot be improved in general. For example, $\zeta(K_{1,n})=1$ and $\operatorname{lcapt}(K_{1,n})=n=\ell(K_{1,n})$.
@@
Let $T$ be a tree as stated. By [18, Theorem 8] we have $\zeta(T)=2$. We use a simplified version of the cop’s strategy provided in the proof of Theorem 2.2, making sure that we remove at least two new leaves from the set of vertices on which the robber might be located in the rest of the game with each move, and removing at least four in the first move. Hence, our strategy will use at most $\lfloor\frac{l(T)}{2}\rfloor-1$ moves, as desired.
@@ 3 Outerplanar graphs
Bonato and Kinnersley showed in [5] that $\zeta(G)\leq 2$ for every outerplanar graph $G$. In this section we extend their results by considering the localization game capture time of outerplanar graphs in the game with two cops. Our results prove that outerplanar graphs with localization number 2 are well-localizable. It remains to consider outerplanar graphs with localization number 1 to fully answer Question 1.2 for all outerplanar graphs.
Recall that a block of a graph is a maximal connected subgraph that has no cut-vertex. The next result relies strongly on the winning strategy given in [5]. For clarity, we only point out the differences and additions made to their results, but for completeness, we recall a few basic definitions.
The two cops playing on an outerplanar graph maintain a cop territory during the game which is a set of vertices with specific properties. Two distinct vertices of the cop territory are the endpoints, sometimes refered to as $v_{L}$ and $v_{R}$. The cop territory is a connected subgraph of the outerplanar graph, after every probe the cops know that the robber is not on a vertex from the cop territory, no vertex (except possibly the endpoints) of the cop territory is adjacent to a vertex outside the cop territory, and both endpoints belong to the same block.
@@
We consider the winning strategy for two cops on an outerplanar graph $G$ described in the proof of [5, Theorem 3.1]. In that strategy, at least one vertex is added to the cop territory in each move, except in the Case 4(b). If $w_{2}=v_{r+1}$, we would have to use two moves to add just one vertex to the cop territory. However, we may easily fix this by skipping the probe on $v_{s}$ and $w_{2}$ and probe instead $v_{L}$ and $w_{2}$ directly, thus saving a move.
@@
Similarly as in the proof of [5, Theorem 3.1] we consider two cases. In each one, we choose vertices $v_{L}$ and $v_{R}$ with $L>R$, where each of them lies on a chord contained in $v_{R}-v_{L}$, such that (after the procedure described in each case is completed) either the robber is captured, or we know the distances $d_{R}$ and $d_{L}$ from the robber to the vertices $v_{R}$ and $v_{L}$ respectively, and the candidate set is contained in $v_{R}-v_{L}$.
@@
If the robber is closer to one endpoint, we do the same as in Case 1. Suppose therefore that $d_{L}=d_{R}$ and let $v_{s}$ be the common neighbor of $v_{R}$ and $v_{L}$ in $v_{R}-v_{L}$. We use a modified version of the argument from Case 4(b) in the proof of [5, Theorem 3.1], discarding at least one chord with each move. Let $v_{R}=w_{1},w_{2},\ldots,w_{k}$ be the neighbors of $v_{s}$ in $G$ that are counterclockwise from $v_{s}$. For $i\in\{1,\ldots,k-1\}$ let sector $i$ be $w_{i}-w_{i+1}$. The cops aim to determine which sector (if any) the robber occupies. We again consider two cases.
@@ 4 Generalization
  • Row 1 consists of all singletons containing a vertex of $G$.
@@
If $S$ is from row $i+1$ and $T$ is from row $i$, then $S$ is a parent of $T$ and $T$ is a child of $S$ if there exists an adjoined coloring of $S$ in which $T$ is monochromatic.
@@
Next, we apply the same procedure on one row lower (with respect to the newly obtained highest but one row) to discard some of the sets. Continuing until we reach row 1 we obtain the desired reduced structure. By definition, every set in row $i+1$ has a child in row $i$. Taking a minimal such set (with respect to inclusion) shows, that the reduced structure has the same amount of rows as the original one, i.e. no rows have become empty during the procedure.
@@
Let $T$ be the tree in Figure 3 and let colorings of $V(T)$ be obtained from comparing distances to a given vertex (see distance colorings below). Then the obtained color classes are
@@
Given a graph $G$ on $n$ vertices and a positive integer $k$, the set of distance colorings is obtained by taking every $S\subseteq V(G)$, $|S|\leq k$, and defining the coloring $C_{S}\colon V(G)\to[n]^{k}$ as $C_{S}(v)=(d(v,s_{1}),\ldots,d(v,s_{\ell}))$ where $S=\{s_{1},\ldots,s_{\ell}\}$.
@@
Let $G$ be a graph, let $k\geq\zeta(G)$ be a positive integer, and let $C_{1},\ldots,C_{m}$ be the distance colorings of $G$. Then the number of rows obtained in this structure is $lcapt_{k}(G)$.
@@
Keeping this in mind, notice also that due to $k\geq\zeta(G)$, every nonempty subset of $V(G)$ appears in the game structure. Row 2 contains all subsets $S$ of $V(G)$ for which there exists a probe on some other vertex set $T$, such that if the candidate set is ever equal to the (closed) neighborhood of $S$, the cops may probe on $T$, corresponding to the coloring $C_{T}$ in which all vertices of the neighborhood of $S$ are colored distinctly (that is, the cops differentiate between these vertices and can thus determine robber’s location). In other words, row 2 contains precisely those vertex sets such that if their neighborhoods are a candidate set, the cops capture the robber in the next round. Analogously, the row $i+1$ contains those vertex sets $S$, whose neighborhoods, should they become a candidate set, can be probed with the probe corresponding to a respective adjoined coloring of $S$, after which the cops know for sure that the robber is located on one of the sets in the lower rows. Hence, inductively, the row $i>1$ contains precisely those vertex sets, such that if their neighborhoods should become candidate sets, the robber can be captured in $i$ moves. Thus clearly the height of the game structure is at most $\operatorname{lcapt}_{k}(G)$.
@@
We similarly provide a strategy for the robber. In the beginning, the candidate set is simply $V(G)$ which is located in the highest row, due to the above considerations. Suppose the candidate set in round $j$ is the neighborhood of a set $S$ in row $i$. Whichever probe the cops choose in round $j+1$, the robber may always make sure that the candidate set is the neighborhood of some set lying on row i1i-1 or higher – if the cops probe on $T$, he chooses as his “position” (he need not choose a single vertex since he knows the moves of the cops in advance) the highest lying color class of $C_{T}(N[S])$. Hence $\operatorname{lcapt}_{k}(G)$ must be greater than or equal to the height of the game structure. ∎
@@
Note that a positive answer to Question 4.4 would also provide an upper bound for $\operatorname{lcapt}(G)$. In particular, proving that $f(n)=Cn$ for some constant $C$ suffices would imply the correctness of Conjecture 1.1.
For example, if we are given only two colorings, we are able to answer Question 4.4 positively.

### v2

# Localization game capture time of trees and outerplanar graphs

@@ Abstract
The localization game is a variant of the game of Cops and Robber in which the robber is invisible and moves between adjacent vertices, but the cops can probe any $k$ vertices of the graph to obtain the distance between probed vertices and the robber. The localization number of a graph is the minimum $k$ needed for the cops to be able to locate the robber in finite time. The localization capture time is the minimum number of rounds needed for the cops to locate the robber.
@@ 1 Introduction
The Cops and Robber game on graphs has been widely studied; see for example [8]. The game is played on a graph by two players, one controlling the cops and the other the robber. Positions of the cops and the robber are known to both players at all times. The players are allowed to move to an adjacent vertex or stay on the same vertex, and players alternate taking moves. The goal of the cops is to capture the robber, i.e. at least one cop must occupy the same vertex as the robber. The robber’s goal is to evade capture forever. The minimum number of cops needed to ensure cops winning on a graph $G$ is the cop number $c(G)$ of $G$. In 2009, Bonato et al. [3] introduced the capture time of a graph as the minimum number of moves needed for $c(G)$ cops to win on $G$. The problem is well-understood for cop-win graphs, but only a handful of results are known for graphs with a larger cop number [16, 17, 13].
@@
In 2012, Seager [18] introduced a version of Cops and Robber where the robber is invisible and the cops are locating the robber using probes. This was later refined to the game called the localization game which has been further studied for example in [4, 5, 6, 7, 10, 11, 12, 13, 14, 15]. This game is also played on a graph by two players, one controlling the cops and the other the robber. However, the robber is invisible to the cops during the game, while the robber knows the positions of the cops. The players alternate taking moves, with the robber selecting the starting vertex first. The robber is allowed to stay on the same vertex or move to a neighbor. The cops are allowed to move to any set of vertices (not necessarily adjacent to their previous positions), and after they all move, each cop sends out a cop probe, which gives the distance between each cop to the robber. Thus after every move of the cops, they obtain a distance vector $(d(c_{1},r),\ldots,d(c_{k},r))$ where $c_{1},\ldots,c_{k}$ are positions of the cops and $r$ is the position of the robber. The goal of the cops is to determine the position of the robber after finitely many rounds. The robber’s goal is to evade capture forever. Additionally, we assume that the robber is omniscient, meaning that the robber is aware of all strategies of the cops. The localization number $\zeta(G)$ of $G$ is the minimum number of cops needed to ensure that the cops win the localization game on $G$. We only consider this game for connected graphs. The exact upper bound for the localization number is known for example for trees, outerplanar graphs [6], and some diameter 2 graphs [4], and is known to be unbounded for planar graphs [10].
@@
In 2022, Behague et al. [1] introduced in full generality the analogous problem to the capture time for the Cops and Robber game for the localization game. The $k$-localization capture time $\operatorname{lcapt}_{k}(G)$ is the minimum number of rounds needed for $k\geq\zeta(G)$ cops to capture the robber on $G$ if the cops are minimizing the number of rounds and the robber is maximizing it. If $k=\zeta(G)$, we refer to the parameter as the localization capture time and denote it by $\operatorname{lcapt}(G)$. Recall that the game starts by the robber choosing an initial vertex. Then, round 1 begins. Each round consists of a move of the cops, after which they obtain a distance vector to the robber’s position (we also say that the cops probe the $k$ vertices they move to), which is followed by a move of the robber. The number of rounds needed for the cops to determine the position of the robber is thus equal to the number of times the cops obtain a distance vector to the robber, i.e. the number of times the cops probe the vertices. Note that as $k$ increases, $\operatorname{lcapt}_{k}(G)$ decreases, and it equals 1 if $k$ is the metric dimension of $G$.
@@ 2 Trees
It is known that one cop wins the localization game on a tree $T$ if and only if $T$ does not contain a copy of $T_{3,3}$ ([19, Theorem 8]). As mentioned in [9] and explicitly proven in [2], it is known that if $T$ is a tree, then $\zeta(T)\leq 2$. It follows from [19] and [1] that the best known upper bound for the localization capture time for a tree $T$ on $n$ vertices is
@@
In this section we improve on these two upper bounds. First observe that $\operatorname{lcapt}(K_{1})=\operatorname{lcapt}(K_{2})=1$ and that $\operatorname{lcapt}(K_{1,m})=m-1=|V(K_{1,m})|-2$ for $m\geq 2$. As $\ell(T)\leq n-2$ for every tree on $n\geq 3$ vertices that is not a star, and $\ell(T)\leq n-4$ for every tree on $n$ vertices that contains a copy of $T_{3,3}$, we get the following result from Theorems 2.2 and 2.3 which are proved below.
@@
If $T$ is a tree on $n\geq 3$ vertices, then
$\operatorname{lcapt}(T)\leq\begin{cases}n-2;&\text{if }\zeta(T)=1,\\ \left\lfloor\frac{n}{2}\right\rfloor-3;&\text{if }\zeta(T)=2.\end{cases}$
As already mentioned, the bound n2n-2 cannot be improved for trees $T$ with $\zeta(T)=1$. After proving Theorem 2.3, we also provide an example showing that the bound is tight for trees $T$ with $\zeta(T)=2$. In the proofs, the candidate set denotes the set of all possible locations of the robber on a graph after a move by the cops or the robber is completed.
@@
If $T$ is a tree on at least two vertices that does not contain a copy of $T_{3,3}$, then $\operatorname{lcapt}(T)\leq\ell(T)$.
@@
Let $T$ be a tree which does not contain a copy of $T_{3,3}$, and root $T$ at any leaf $v_{0}$. We provide a strategy for the cop to construct a sequence of pairwise distinct vertices $v_{1},v_{2},\ldots,v_{a}$ of $T$, lying on levels $k_{1}<k_{2}<\cdots<k_{a}$ respectively, such that for each $1\leq i\leq a$, the candidate set will be contained in the set of descendants of $v_{i}$, immediately after the cop’s probe. In particular, after each application of one of the strategies described below, the cop will be able to either capture the robber, or identify the lowest positioned vertex $v_{i}$, on level $k_{i}>k_{i-1}$, such that the candidate set will be contained in the set of its descendants throughout the remainder of the game. Since the tree is finite, the cop will indeed capture the robber in a finite number of moves. Moreover, we prove that during the described strategy, the cop (in most cases) maintains the following property: the number of probes needed to eliminate a part of the tree from all the possible future candidate sets with respect to the outlined strategy is at most the number of leaves in that part of the tree. There are a few exceptions to this general rule which are dealt with on the spot.
Choose $v_{0}$ as the first probe. If the robber is not captured, then the candidate set is clearly contained in the set of descendants of $v_{0}$. Hence, there exists a unique vertex $v$, positioned lowest out of all the vertices for which the candidate set is contained in the set of their descendants. We set $v_{1}=v$. Suppose that after some move, the cop has already constructed $v_{1},\ldots,v_{t}$. Since the robber must be located on a descendant of $v_{t}$, we may now consider the levels of these vertices with respect to $v_{t}$, that is, as if the tree was rooted at $v_{t}$, in order to simplify the notation. Recall that the vertex set consisting of a child of $v_{t}$ and all of its descendants is called a branch.
@@
Suppose now that there is at least one child of $v_{t}$ of degree 2. Again, if such children are not present, the cop simply omits this step. Otherwise, he consecutively probes the children of such vertices. Let $u$ be one such vertex that the cop probes, and let $d$ be the answer he gets after this probe. Immediately before the probe, the robber cannot be on any other level than k1,kk-1,k or $k+1$. We consider the following four cases. In all of them, if the cop ever realizes that the robber is on the same branch as him and on level 3 or lower, we may simply set $v_{t+1}$ to be $u$ or, if possible, some other vertex on the branch on a lower level.
@@
Notice that in each case, the cop has eliminated some branches as the robber’s possible whereabouts, using at most as many moves as there are leaves on the discarded branches. Hence, using the described strategy, the cop extends his sequence of vertices, until the robber is captured, needing at most $\ell(T)$ moves to do so. ∎
We do not know if the upper bound in Theorem 2.2 is tight. However, many trees $T$, including paths and stars, have capture time equal to $\ell(T)-1$.
@@
Let $T$ be a tree as stated. By [19, Theorem 8] we have $\zeta(T)=2$. We use a simplified version of the cop’s strategy provided in the proof of Theorem 2.2, making sure that we remove at least two new leaves from the set of vertices on which the robber might be located in the rest of the game with each move, and removing at least four in the first move. Hence, our strategy will use at most $\lfloor\frac{\ell(T)}{2}\rfloor-1$ moves, as desired.
@@ 3 Outerplanar graphs
Bonato and Kinnersley showed in [6] that $\zeta(G)\leq 2$ for every outerplanar graph $G$. In this section we extend their results by considering the localization capture time of outerplanar graphs in the game with two cops. Our results prove that outerplanar graphs with localization number 2 are well-localizable. It remains to consider outerplanar graphs with localization number 1 to fully answer Question 1.2 for all outerplanar graphs.
Recall that a block of a graph is a maximal connected subgraph that has no cut-vertex. The next result relies strongly on the winning strategy given in [6]. For clarity, we only point out the differences and additions made to their results. We recall a few basic definitions.
The two cops playing on an outerplanar graph maintain a cop territory during the game which is a set of vertices with specific properties. Two distinct vertices of the cop territory are the endpoints, sometimes referred to as $v_{L}$ and $v_{R}$. The cop territory is a connected subgraph of the outerplanar graph, after every probe the cops know that the robber is not on a vertex from the cop territory, no vertex (except possibly the endpoints) of the cop territory is adjacent to a vertex outside the cop territory, and both endpoints belong to the same block.
@@
We consider the winning strategy for two cops on an outerplanar graph $G$ described in the proof of [6, Theorem 3.1]. In that strategy, at least one vertex is added to the cop territory in each move, except in the Case 4(b). If $w_{2}=v_{r+1}$, we would have to use two moves to add just one vertex to the cop territory. However, we may easily fix this by skipping the probe on $v_{s}$ and $w_{2}$ and probe instead $v_{L}$ and $w_{2}$ directly, thus saving a move.
@@
Similarly as in the proof of [6, Theorem 3.1] we consider two cases. In each one, we choose vertices $v_{L}$ and $v_{R}$ with $L>R$, where each of them lies on a chord contained in $v_{R}-v_{L}$, such that (after the procedure described in each case is completed) either the robber is captured, or we know the distances $d_{R}$ and $d_{L}$ from the robber to the vertices $v_{R}$ and $v_{L}$ respectively, and the candidate set is contained in $v_{R}-v_{L}$.
@@
If the robber is closer to one endpoint, we do the same as in Case 1. Suppose therefore that $d_{L}=d_{R}$ and let $v_{s}$ be the common neighbor of $v_{R}$ and $v_{L}$ in $v_{R}-v_{L}$. We use a modified version of the argument from Case 4(b) in the proof of [6, Theorem 3.1], discarding at least one chord with each move. Let $v_{R}=w_{1},w_{2},\ldots,w_{k}$ be the neighbors of $v_{s}$ in $G$ that are counterclockwise from $v_{s}$. For $i\in\{1,\ldots,k-1\}$ let sector $i$ be $w_{i}-w_{i+1}$. The cops aim to determine which sector (if any) the robber occupies. We again consider two cases.
@@ 4 Generalization
For $S\subseteq V(G)$ and a coloring $C$ of $V(G)$, we denote by $C(S)$ the restriction of $C$ to $S$ which is also a coloring. Recall that a color class of a coloring is a subset of vertices that are assigned the same color.
@@
  • Row 1 consists of singletons $\{v\}$ for every $v\in V(G)$.
@@
If $S$ is in row $i+1$ and $T$ is in row $i$, then $S$ is a parent of $T$ and $T$ is a child of $S$ if there exists a coloring adjoined to the set $S$, in which $T$ is monochromatic.
@@
Next, we apply the same procedure on one row lower (with respect to the newly obtained highest but one row) to discard some of the sets. Continuing until we reach row 1 we obtain the desired reduced structure. By definition, every set in row $i+1$ has a child in row $i$. Taking a minimal such set (with respect to inclusion) shows that the reduced structure has the same amount of rows as the original one, i.e. no rows have become empty during the procedure.
@@
Let $T$ be the tree in Figure 3 and let colorings of $V(T)$ be obtained from comparing distances to a given vertex. Then the obtained color classes are
@@
Given a graph $G$ on $n$ vertices and a positive integer $k$, the set of distance colorings is obtained by taking every $S\subseteq V(G)$, $|S|=k$, and defining the coloring $C_{S}\colon V(G)\to[n]^{k}$ as $C_{S}(v)=(d(v,s_{1}),\ldots,d(v,s_{\ell}))$ where $S=\{s_{1},\ldots,s_{\ell}\}$. Note that this set of colorings is exactly the one we used in Example 4.1 with $k=1$.
@@
Let $G$ be a graph, let $k\geq\zeta(G)$ be a positive integer, and let $C_{1},\ldots,C_{m}$ be the distance colorings of $G$. Then the number of rows obtained in this structure is $lcapt_{k}(G)+1$.
@@
Keeping this in mind, notice also that due to $k\geq\zeta(G)$, every nonempty subset of $V(G)$ appears in the game structure. Row 2 contains all subsets $S$ of $V(G)$ for which there exists a probe on some other vertex set $T$, such that if the candidate set is ever equal to the (closed) neighborhood of $S$, the cops may probe on $T$, corresponding to the coloring $C_{T}$ in which all vertices of the neighborhood of $S$ are colored distinctly (that is, the cops differentiate between these vertices and can thus determine robber’s location). In other words, row 2 contains precisely those vertex sets such that if their neighborhoods are a candidate set, the cops capture the robber in the next round. Analogously, the row $i+1$ contains those vertex sets $S$, whose neighborhoods, should they become a candidate set, can be probed with the probe corresponding to a respective adjoined coloring of $S$, after which the cops know for sure that the robber is located on one of the sets in the lower rows. Hence, inductively, the row $i>1$ contains precisely those vertex sets, such that if their neighborhoods should become candidate sets, the robber can be captured in i1i-1 moves. Thus clearly the height of the game structure is at most $\operatorname{lcapt}_{k}(G)+1$.
@@
We similarly provide a strategy for the robber. In the beginning, the candidate set is simply $V(G)$ which is located in the highest row, due to the above considerations. Suppose the candidate set in round $j$ is the neighborhood of a set $S$ in row $i$. Whichever probe the cops choose in round $j+1$, the robber may always make sure that the candidate set is the neighborhood of some set lying on row i1i-1 or higher – if the cops probe on $T$, he chooses as his “position” (he need not choose a single vertex since he knows the moves of the cops in advance) the highest lying color class of $C_{T}(N[S])$. Hence $\operatorname{lcapt}_{k}(G)+1$ must be greater than or equal to the height of the game structure (the robber is already captured on row 1). ∎
If we wish to also include colorings corresponding to the cops probing on vertices which are not necessarily distinct, we can relax the condition $|S|=k$ to $|S|\leq k$. But as long as $k\leq|G|$, the cops need never choose such a probe in their optimal strategy, hence such colorings are irrelevant for our discussion.
Example 4.3.
Let $T$ be the tree $T_{3,3}$ in Figure 4 and let colorings of $V(T)$ be distance colorings obtained from two cops.
0112233445566778899
Figure 4: Tree $T_{3,3}$ with the distance coloring on vertices $1$ and $2$.
The game structure is:
Row 3: Any set of vertices whose closed neighborhood contains
$\text{at least five of the vertices $4$, $5$, $6$, $7$, $8$, $9$};$
[…]
Row 1: $0,1,2,3,4,5,6,7,8,9.$
To see how the game structure simulates the game, let the robber choose the set of all vertices of the tree as it’s starting vertex set (which is of course contained in the topmost row). This represents the starting position of the game (the candidate set is the entire graph), when the robber chooses an initial vertex, say vertex $4$. Then there exists a coloring of $T_{3,3}$ with color classes lying in rows $1$ and $2$. This is the optimal coloring (representing a probe) that the cops can choose in the game. In our case, one such coloring is the distance coloring on vertices $1$ and $2$, as shown in Figure 4. The robber can now choose any of the color classes of this coloring, and assuming optimal play, he chooses one lying in row $2$, for example $45$. This corresponds to the robber staying on vertex $4$. In the next round, the candidate set is $145$, and the cops capture the robber by choosing the distance coloring on vertices $4$ and $5$, which, of course, corresponds to a probe on vertices $4$ and $5$ in the actual game. In this way, every possible game is modeled in the game structure.
The reduced game structure is then:
Row 3: $123,128,129,136,137,234,235,1679,1678,1689,1789,2458,2459,2489,2589,$
$3456,3457,3467,3567,45678,45679,45689,45789,46789,56789;$
Row 2: $04,05,06,07,08,09,45,67,89,12,13,23;$
Row 1: $0,1,2,3,4,5,6,7,8,9.$
This reduces the number of sets in the structure, while retaining its height and the relevant parent-child relationships.
@@
Note that a positive answer to Question 4.5 would also provide an upper bound for $\operatorname{lcapt}(G)$. In particular, proving that $f(n)=Cn$ for some constant $C$ suffices would imply the correctness of Conjecture 1.1.
For example, if we are given only two colorings, we are able to answer Question 4.5 positively.